# Controllable Reference-Based Real-World Remote Sensing Image Super-Resolution with Generative Diffusion Priors

### v1

@@ Abstract
Super-resolution (SR) techniques can enhance the spatial resolution of remote sensing images by utilizing low-resolution (LR) images to reconstruct high-resolution (HR) images, enabling more efficient large-scale earth observation applications. While single-image super-resolution (SISR) methods have shown progress, reference-based super-resolution (RefSR) offers superior performance by incorporating historical HR images alongside current LR observations. However, existing RefSR methods struggle with real-world complexities, such as cross-sensor resolution gap and significant land cover changes, often leading to under-generation or over-reliance on reference image. To address these challenges, we propose CRefDiff, a novel controllable reference-based diffusion model for real-world remote sensing image SR. To address the under-generation problem, CRefDiff is built upon the pretrained Stable Diffusion model, leveraging its powerful generative prior to produce accurate structures and textures. To mitigate over-reliance on the reference, we introduce a dual-branch fusion mechanism that adaptively integrates both local and global information from the reference image. Moreover, this novel dual-branch design enables reference strength control during inference, enhancing the model’s interactivity and flexibility. Finally, a strategy named Better Start is proposed to significantly reduce the number of denoising steps, thereby accelerating the inference process. To support further research, we introduce Real-RefRSSRD, a new real-world RefSR dataset for remote sensing images, consisting of HR NAIP and LR Sentinel-2 image pairs with diverse land cover changes and significant temporal gaps. Extensive experiments on Real-RefRSSRD show that CRefDiff achieves state-of-the-art performance across various metrics and improves downstream tasks such as scene classification and semantic segmentation. The dataset and code will be publicly available at https://github.com/wwangcece/CRefDiff
@@ 1 Introduction
However, when dealing with large upscaling factors, the aforementioned generative approaches often suffer from uncontrollable hallucination issues [93], where the model produces unrealistic or non-existent textures during the SR process. To mitigate this issue, recent studies have begun to shift from single-image SR (SISR) to reference-based SR (RefSR). By incorporating additional HR reference images, RefSR enables the reconstruction of high-frequency details with improved fidelity. For natural images, the reference is usually obtained from different viewpoints of the same object [94, 82] or is retrieved using web search engines [24]. In the context of remote sensing imagery, Dong et al. [10] proposed using HR images from platforms such as Google Earth as references and constructed the first RefSR dataset for remote sensing imagery. However, these datasets generate LR images using bicubic downsampling from corresponding HR images, without accounting for the complex degradations introduced by cross-sensor differences in real-world scenarios. As a result, its applicability to practical remote sensing tasks is limited. To overcome the aforementioned limitations, we propose leveraging historical HR observations together with current LR observations to generate the current HR observation, as formulated below:
@@
Here, $I_{\mathrm{ref}}$ and $I_{\mathrm{hr}}$ denote the historical and current observations from a high-resolution satellite, while $I_{\mathrm{lr}}$ represents the current observation from a low-resolution satellite. Therefore, once trained, the model $f_{\theta}$ can predict the current HR observations based on a historical HR observation and frequently acquired LR observations.
@@
To address the aforementioned challenges, we propose a Controllable Reference-based Diffusion Model (CRefDiff) tailored for the real-world RefSR task. Built upon the pretrained Stable Diffusion (SD) model [52], CRefDiff leverages its strong generative capacity to bridge the substantial resolution gap (exceeding 8×8$\times$8 ×) between different satellite platforms. To more effectively exploit both local and global information from the reference image, we introduce a Change Aware Attention (CAA) block and a Semantic Token Aggregation (STA) module, enabling the model to incorporate reference cues at both local and global levels. To alleviate the heavy computational burden caused by the iterative sampling process of diffusion models, we design a Better Start strategy that modifies the starting point of the reverse process, effectively shortening the denoising trajectory. To fundamentally mitigate the model’s over-reliance on the reference image, we further introduce a reference strength control mechanism during inference. This mechanism allows users to adaptively adjust the degree to which the model leverages reference information, thereby enhancing both the interactivity and flexibility of the model. The main contributions of this work are summarized as follows:
@@
    We propose CRefDiff, a novel controllable reference-based super-resolution framework built on pretrained generative diffusion models. It effectively bridges the resolution gap between cross-sensor inputs and mitigates hallucinations through a carefully designed local textures and global contextual fusion architecture tailored for remote sensing imagery.
@@
    We design two novel inference-time mechanisms, Better Start and Reference Strength Control. Better Start shortens the diffusion trajectory, significantly accelerating inference. While, Reference Strength Control allows users to dynamically modulate the influence of the reference image both globally and spatially, enhancing flexibility and robustness under different changing conditions.
@@ 2.1 Remote Sensing Image Super-Resolution
In recent years, with the rapid advancement of deep learning techniques, numerous deep learning-based SR methods have been proposed, significantly outperforming traditional approaches [59, 83, 61, 13]. Dong et al. [9] were the first to introduce neural networks into the SR task. Since then, extensive efforts have been devoted to improving SR performance by designing more effective network architectures, such as residual blocks [30], dense blocks [28], and recursive blocks [29]. Compared with natural images, remote sensing imagery possesses unique characteristics. Consequently, numerous studies have adapted network architectures to better suit these properties, such as multi-scale networks [31, 5, 32], multi-stage architectures [33, 37], gradient guidance mechanisms [23, 49], and scene-adaptive networks [35, 91].
@@
Recently, denoising diffusion probabilistic models [18] (DDPMs) have demonstrated remarkable advantages in the field of AI-generated content (AIGC). Compared to GANs or normalizing flow models, DDPMs offer more stable training, greater architectural flexibility, and better scalability [84]. As a result, an increasing number of researchers have begun to explore the application of DDPMs to the SR task. Saharia et al. [53] were the first to employ the DDPM architecture for image SR, achieving performance comparable to that of GAN-based methods. Yue et al. [85] introduced the concept of residual diffusion to better exploit the available pixels in LR images, significantly reducing inference time. Xiao et al. [77] were the first to apply diffusion models to remote sensing image SR and proposed an efficient denoising network tailored for this task. To further enhance the robustness and generalization of SR models in unseen scenarios, an increasing number of studies have begun to develop SR methods based on large-scale text-to-image pretrained models, such as Stable Diffusion [52]. Wang et al. [70] were the first to explore this direction, significantly improving the perceptual quality of reconstruction results. Khanna et al. [27] fine-tuned a pretrained SD model using remote sensing imagery and associated metadata, achieving 10×10$\times$10 × SR for Sentinel-2 images.
Although generative models alleviate over-smoothing issues, they can also introduce hallucinated content, generating unrealistic textures or structures. To address these limitations, the following section introduces another SR paradigm—reference-based super-resolution.
@@ 3 Preliminary
3.1 Naive Solution and Problems
Unlike natural images, reference satellite images are usually coarsely aligned with HR images in the spatial domain. Therefore, to fuse the information from a roughly aligned reference image and a LR image, a naive solution is to simply concatenate LR and reference images, or features of them, and feed the result into a SR backbone model (in our case, ESRGAN [72]) for end-to-end training. However, we observe that this approach struggles to handle regions with significant land cover changes. As illustrated in Fig. 2, when new objects appear in the target HR image, the model, relying primarily on the LR input, fails to reconstruct accurate structures, a phenomenon we refer to as under-generation. Conversely, when certain objects present in the reference image are missing in the HR image, the model tends to over-rely on the reference and hallucinates incorrect content, which we term the over-reliance problem. As shown in Fig. 3, we present three similarity metrics between the ground truth (GT) image and the reference and LR images. Despite the presence of land cover changes, the reference image consistently exhibits a higher similarity to the GT image across all metrics. This demonstrates that the changed regions between the reference and HR images are notably fewer than the unchanged ones, which leads the network to favor a direct “copy-and-paste” strategy from the reference image [81]. This indicates that the naive concatenation strategy hinders the network from effective learning the underlying changes between the LR and the reference images, thereby limiting its ability to handle regions with significant land cover changes during inference.
Figure 2: In the presence of substantial land cover changes, the naive solution often exhibits issues of under-generation (failing to recover new structures) and over-reliance on the reference (hallucinating outdated or incorrect content).
@@ 4.1 Overview
During inference, we introduce two optional strategies: Better Start and Reference Strength Control. The Better Start strategy leverages the output of a naive solution as the initial state for the reverse diffusion process. This significantly shortens the denoising trajectory and reduces the number of required inference steps. The Reference Strength Control strategy enables users to specify either a scalar $s$ or a spatial mask $m$ to control the degree to which reference information influences the model, at both global and local levels. This design fundamentally mitigates the over-reliance problem by allowing adaptive regulation of reference usage.
@@ 4.2.1 Local Texture Encoder
Although local cross-attention can adaptively extract useful information within each window, it treats all windows equally and thus fails to handle large-scale land cover changes—especially those beyond the window size. To address this, we introduce a similarity mask into the residual connection of the cross-attention module to regulate the degree of reference information integration:
@@ 4.2.2 Adapter Design
In this way, the information from the LR and reference images can be effectively aligned with the inherent priors of the pre-trained SD model, thereby guiding its denoising process.
@@ 4.5 Optimization and Inference
This enables us to replace $z_{T}$ with $z_{t^{\prime}}^{\mathrm{start}}$ as the initial state of the reverse denoising process, thereby reducing the number of required denoising steps from $T$ to $t^{\prime}$. However, for real-world satellite imagery, the LR image often differs significantly from the GT image in both spatial resolution and visual style. Directly using the LR image as the starting point significantly degrades the inference quality. Therefore, we adopt the output of an off-the-shelf SR model [72] as the starting point, which effectively reduces the number of inference steps while maintaining satisfactory performance.
@@
For the global branch, this can be achieved by modifying the cross-attention mechanism. Specifically, we use the factor $s$ to modulate the residual connection of the cross-attention, and apply the mask $m$ as the attention mask to control the influence of the reference information:
@@ 5.5.1 Effectiveness of Global Branch
Local Global SD prior PSNR\ $\uparrow$ LPIPS\ $\downarrow$ CLIPIQA\ $\uparrow$
$\checkmark$ $\times$ $\checkmark$ 20.86 0.3926 0.5855
$\times$ $\checkmark$ $\checkmark$ 21.18 0.4006 0.5909
$\checkmark$ $\times$ $\times$ 22.13 0.6280 0.1533
$\checkmark$ $\checkmark$ $\checkmark$ 21.33 0.3852 0.6014
@@
PSNR\ $\uparrow$ LPIPS\ $\downarrow$ CLIPIQA\ $\uparrow$
CLIP 20.97 0.3917 0.5950
MLP 21.37 0.3885 0.5994
Ours 21.33 0.3852 0.6014

### v2

# Controllable Reference-Guided Diffusion with Local–Global Fusion for Real-World Remote Sensing Image Super-Resolution

@@ Abstract
Super-resolution (SR) techniques can enhance the spatial resolution of remote sensing images, enabling more efficient large-scale earth observation applications. While single-image super-resolution methods enhance low-resolution (LR) images, they neglect valuable complementary information from auxiliary data. Reference-based super-resolution (RefSR) can be interpreted as an information fusion task, where historical high-resolution (HR) reference images are combined with current LR observations. However, existing RefSR methods struggle with real-world complexities, such as cross-sensor resolution gap and significant land cover changes, often leading to under-generation or over-reliance on reference image. To address these challenges, we formulate the RefSR task as a multi-sensor fusion problem and propose CRefDiff, a novel controllable reference-guided diffusion model for real-world remote sensing image super-resolution. To address the under-generation problem, CRefDiff leverages a powerful generative prior to produce accurate structures and textures. To mitigate over-reliance on the reference, we introduce a dual-branch fusion mechanism that adaptively fuse both local and global information from the reference image. Moreover, this novel dual-branch design enables reference strength control during inference, enhancing the model’s interactivity and flexibility. Finally, a strategy named Better Start is proposed to significantly reduce the number of denoising steps, thereby accelerating the inference process. To support further research, we introduce Real-RefRSSRD, a new real-world RefSR dataset for remote sensing images, consisting of HR NAIP and LR Sentinel-2 image pairs with diverse land cover changes and significant temporal gaps. Extensive experiments on Real-RefRSSRD show that CRefDiff achieves state-of-the-art performance across various metrics and improves downstream tasks such as scene classification and semantic segmentation. The dataset and code is publicly available at https://github.com/wwangcece/CRefDiff
@@ 1 Introduction
However, when dealing with large upscaling factors, the aforementioned generative approaches often suffer from uncontrollable hallucination issues [95], where the model produces unrealistic or non-existent textures during the SR process. To mitigate this issue, recent studies have begun to shift from single-image SR (SISR) to reference-based SR (RefSR). By incorporating additional HR reference images, RefSR enables the reconstruction of high-frequency details with improved fidelity. For natural images, the reference is usually obtained from different viewpoints of the same object [96, 84] or is retrieved using web search engines [24]. In the context of remote sensing imagery, Dong et al. [10] proposed using HR images from platforms such as Google Earth as references and constructed the first RefSR dataset for remote sensing imagery. However, these datasets generate LR images using bicubic downsampling from corresponding HR images, without accounting for the complex degradations introduced by cross-sensor differences in real-world scenarios. In contrast, we formulate the RefSR task as a fusion problem of heterogeneous remote sensing imagery, where the historical reference images, current LR observations, and current HR targets are all derived from real remote sensing observation:
@@
Here, $I_{\mathrm{ref}}$ and $I_{\mathrm{hr}}$ denote the historical and current observations from a high-resolution satellite, while $I_{\mathrm{lr}}$ represents the current observation from a low-resolution satellite. Therefore, once trained, the model $f_{\theta}$ can generate the current HR observation by integrating a historical HR image with frequently acquired LR observations, thereby enabling high spatiotemporal resolution Earth observation.
@@
To address the aforementioned challenges, we propose a Controllable Reference-Guided Diffusion (CRefDiff) model tailored for the real-world RefSR task. Built upon the pretrained Stable Diffusion (SD) model [53], CRefDiff leverages its strong generative capacity to bridge the substantial resolution gap (exceeding $8\times$) between different satellite platforms. To more effectively fuse both local and global information from the reference image, we introduce a Change Aware Attention (CAA) block and a Semantic Token Aggregation (STA) module, enabling the model to incorporate reference cues at both local and global levels. To alleviate the heavy computational burden caused by the iterative sampling process of diffusion models, we design a Better Start strategy that modifies the starting point of the reverse process, effectively shortening the denoising trajectory. To fundamentally mitigate the model’s over-reliance on the reference image, we further introduce a reference strength control mechanism during inference. This mechanism allows users to adaptively adjust the degree to which the model fuses reference information, thereby enhancing both the interactivity and flexibility of the model. The main contributions of this work are summarized as follows:
@@
    We propose CRefDiff, a novel controllable RefSR framework built on pretrained generative diffusion models. It effectively fuses multi-sensor heterogeneous information and mitigates hallucinations through a carefully designed fusion architecture that jointly models local textures and global contextual cues.
@@
    We design two novel inference-time mechanisms, Better Start and Reference Strength Control. Better Start shortens the diffusion trajectory, significantly accelerating inference. While, Reference Strength Control allows users to dynamically modulate the influence of the reference image, enhancing fusion flexibility and robustness under different changing conditions.
@@ 2.1 Remote Sensing Image Super-Resolution
In recent years, with the rapid advancement of deep learning techniques, numerous deep learning-based SR methods have been proposed, significantly outperforming traditional approaches [60, 85, 62, 13]. Dong et al. [9] were the first to introduce neural networks into the SR task. Since then, extensive efforts have been devoted to improving SR performance by designing more effective network architectures, such as residual blocks [30], dense blocks [28], and recursive blocks [29]. Compared with natural images, remote sensing imagery possesses unique characteristic [41, 98]. Consequently, numerous studies have adapted network architectures to better suit these properties, such as multi-scale networks [31, 5, 32], multi-stage architectures [33, 37], gradient guidance mechanisms [23, 50], scene-adaptive networks [35, 93] and degradation-adaptive networks [79].
@@
Recently, denoising diffusion probabilistic models [18] (DDPMs) have demonstrated remarkable advantages in the field of AI-generated content (AIGC). Compared to GANs or normalizing flow models, DDPMs offer more stable training, greater architectural flexibility, and better scalability [86]. As a result, an increasing number of researchers have begun to explore the application of DDPMs to the SR task. Saharia et al. [54] were the first to employ the DDPM architecture for image SR, achieving performance comparable to that of GAN-based methods. Yue et al. [87] introduced the concept of residual diffusion to better exploit the available pixels in LR images, significantly reducing inference time. Xiao et al. [78] were the first to apply diffusion models to remote sensing image SR and proposed an efficient denoising network tailored for this task. To further enhance the robustness and generalization of SR models in unseen scenarios, an increasing number of studies have begun to develop SR methods based on large-scale text-to-image pretrained models, such as Stable Diffusion [53]. Wang et al. [71] were the first to explore this direction, significantly improving the perceptual quality of reconstruction results. Khanna et al. [27] fine-tuned a pretrained SD model using remote sensing imagery and associated metadata, achieving $10\times$ SR for Sentinel-2 images.
Although generative models alleviate over-smoothing issues, they can also introduce hallucinated content, generating unrealistic textures or structures. Unlike previous approaches, our method incorporates reference images and fully exploits both their local fine-grained details and global contextual information. Moreover, it allows users to explicitly control the proportion of information derived from the LR and reference images, thereby enabling more flexible and controllable content generation.
@@ 3 Preliminary
3.1 Problem Statement
Unlike natural images, reference satellite images are usually coarsely aligned with HR images in the spatial domain. Therefore, to fuse the information from a roughly aligned reference image and a LR image, we train and test existing RefSR methods for remote sensing imagery (in our case, RRSGAN [10]) and RGTGAN [63]) on the collected remote sensing imagery. However, we observe these approaches struggle to handle regions with significant land cover changes. As illustrated in Fig. 2, when new objects appear in the target HR image, the model, relying primarily on the LR input, fails to reconstruct accurate structures, a phenomenon we refer to as under-generation. Conversely, when certain objects present in the reference image are missing in the HR image, the model tends to over-rely on the reference and hallucinates incorrect content, which we term the over-reliance problem. As shown in Fig. 3, we present three similarity metrics between the ground truth (GT) image and the reference and LR images. Despite the presence of land cover changes, the reference image consistently exhibits a higher similarity to the GT image across all metrics. This demonstrates that the changed regions between the reference and HR images are notably fewer than the unchanged ones, which leads the network to favor a direct “copy-and-paste” strategy from the reference image [83]. This suggests that existing methods fail to capture the underlying discrepancies between LR and reference images, thereby limiting their ability to handle regions with substantial land-cover changes during inference.
Figure 2: In the presence of substantial land cover changes, current RefSR methods often exhibit issues of under-generation (failing to recover new structures) and over-reliance on the reference (hallucinating outdated or incorrect content).
@@ 4.1 Overview
During inference, we introduce two optional strategies: Better Start and Reference Strength Control. The Better Start strategy leverages the output of a lightweight fusion model as the initial state for the reverse diffusion process. This significantly shortens the denoising trajectory and reduces the number of required inference steps. The Reference Strength Control strategy enables users to specify either a scalar $s$ or a spatial mask $m$ to control the degree to which reference information influences the model, at both global and local levels. This design fundamentally mitigates the over-reliance problem by allowing adaptive regulation of reference usage.
@@ 4.2.1 Local Texture Encoder
Although local cross-attention can adaptively extract useful information within each window, it treats all windows equally and thus fails to handle large-scale land cover changes—especially those beyond the window size. To address this, we first compute the cosine similarity between the features of the LR image and the reference image, and then use it to modulate the residual connections in the cross-attention module:
@@ 4.2.2 Adapter Design
Here, $F_{\mathrm{enc}}^{i}$ and $F_{\mathrm{dec}}^{i}$ represent the features of the denoising UNet’s encoding and decoding layers, respectively. In this way, the information from the LR and reference images can be effectively aligned with the inherent priors of the pre-trained SD model, thereby guiding its denoising process.
@@ 4.5 Optimization and Inference
This enables us to replace $z_{T}$ with $z_{t^{\prime}}^{\mathrm{start}}$ as the initial state of the reverse denoising process, thereby reducing the number of required denoising steps from $T$ to $t^{\prime}$. However, for real-world satellite imagery, the LR image often differs significantly from the GT image in both spatial resolution and visual style. Directly using the LR image as the starting point significantly degrades the inference quality. Therefore, we train a lightweight SR model and perform an initial fusion of the LR and reference images. The resulting preliminary fusion is then used as the starting point for denoising, which effectively reduces the number of inference steps while maintaining satisfactory performance.
@@
For the global branch, this can be achieved by modifying the cross-attention mechanism. Specifically, we use the factor $s$ to modulate the residual connection of the cross-attention $f_{\mathrm{cross\text{-}attn}}$, and apply the mask $m$ as the attention mask to control the influence of the reference information:
@@ 5.5.1 Effectiveness of Global Branch
Local Global SD prior PSNR\ $\uparrow$ SSIM\ $\uparrow$ LPIPS\ $\downarrow$ DISTS\ $\downarrow$ MUSIQ\ $\uparrow$ CLIPIQA\ $\uparrow$
$\checkmark$ $\times$ $\checkmark$ 20.86 0.4564 0.3926 0.2362 52.89 0.5855
$\times$ $\checkmark$ $\checkmark$ 21.18 0.4615 0.4006 0.2414 53.36 0.5909
$\checkmark$ $\times$ $\times$ 22.13 0.5092 0.6280 0.3739 23.94 0.1533
$\checkmark$ $\checkmark$ $\checkmark$ 21.33 0.4691 0.3852 0.2304 54.44 0.6014
@@
PSNR\ $\uparrow$ SSIM\ $\uparrow$ LPIPS\ $\downarrow$ DISTS\ $\downarrow$ MUSIQ\ $\uparrow$ CLIPIQA\ $\uparrow$
CLIP 20.97 0.4557 0.3917 0.2320 53.94 0.5950
MLP 21.37 0.4645 0.3885 0.2371 54.27 0.5994
Ours 21.33 0.4691 0.3852 0.2304 54.44 0.6014
@@ 5.5.3 Effectiveness of SD Prior
5.5.4 Effectiveness of Augmentation Strategy
In this section, we validate the effectiveness of the employed reference image augmentation strategy. This strategy perturbs the style of reference images during training, thereby encouraging the model to focus more on structural content information. Fig. 17 presents reconstruction results without applying this strategy during training. It can be observed that, in this case, the model tends to overly depend on reference information, transferring non-existent objects into the reconstructed images. In addition, the quantitative results in Table 5 also demonstrate the performance gains achieved by this strategy. These findings collectively validate the effectiveness of the proposed strategy.
Figure 17: Visual comparison with and without the reference image augmentation strategy.
Table 5: Quantitative comparison with and without the reference image augmentation strategy.
PSNR\ $\uparrow$ SSIM\ $\uparrow$ LPIPS\ $\downarrow$ DISTS\ $\downarrow$ MUSIQ\ $\uparrow$ CLIPIQA\ $\uparrow$
W/O Augment 21.06 0.4581 0.3951 0.2371 54.14 0.5978
Ours 21.33 0.4691 0.3852 0.2304 54.44 0.6014
@@ 5.6.1 High Spatiotemporal Resolution Image Synthesis
5.6.2 Cross-Dataset Transfer Evaluation
To evaluate the generalization capability of our model, we collected an additional cross-domain dataset for validation. In this dataset, both the HR images and the reference images are obtained from the National Ecological Observatory Network (NEON) with a spatial resolution of 1 m, while the LR images are acquired from Sentinel-2. The dataset was collected in the Edgewater region of the United States, which is not covered by our original dataset in either time or space. We directly evaluate the model trained on our proposed dataset on this cross-domain dataset. The results of our method and four representative RefSR baselines are presented in Table 6. As shown, our method consistently outperforms the comparison approaches, demonstrating superior performance even under cross-domain conditions. This indicates that our approach achieves stronger robustness and generalization in real-world scenarios.
Table 6: Cross-dataset transfer evaluation results. The bold and underlined numbers denote the best and second-best results.
LPIPS\ $\downarrow$ DISTS\ $\downarrow$ MUSIQ\ $\uparrow$ CLIPIQA\ $\uparrow$
TTSR 0.5319 0.3420 57.83 0.6075
DATSR 0.6128 0.3464 37.35 0.4617
RRSGAN 0.5501 0.3372 43.59 0.5957
RGTGAN 0.4734 0.2833 55.13 0.6389
Ours 0.4203 0.2361 63.01 0.7424
Figure 20: Visual comparison of reconstruction results between our method and other RefSR approaches under extreme land-cover changes.
@@ 5.6.3 Effect of Reference Similarity
Fig. 20 presents the visual results of our method and all compared RefSR in scenarios with substantial land-cover changes. In the three representative examples, numerous new buildings appear in the HR images compared to the reference images. The dense and fine-grained structures of these buildings pose significant challenges for accurate reconstruction. Under such conditions, the baseline methods tend to produce either overly blurry or heavily distorted results. In contrast, our approach is able to effectively exploit the reference information and accurately recover the structural details of all objects.
5.6.4 Local-Global Reference Strength Control
Although our proposed CRefDiff demonstrates a strong ability to identify land cover changes between the LR and reference images, some failure cases involving incorrect reference usage still occur under certain challenging scenarios. To address this issue, we propose a Reference Strength Control strategy, which allows users to manually correct such cases. During inference, users can either supply a global scalar factor $s$ to adjust the overall reliance on reference information, or provide a spatial mask $m$ to modulate reference strength locally. As illustrated in Fig. 22, a representative example is shown. In the highlighted region, the HR and reference images exhibit significant structural differences. However, the model still incorporates some misleading reference cues, resulting in erroneous structures in the reconstructed output (as indicated by “s=1s=1”). To alleviate this, the user can progressively reduce the global strength factor to suppress reference guidance. Alternatively, a binary mask specifying unreliable regions can be provided to locally weaken the influence of reference features in those areas.